# Beauty and the Bias: Exploring the Impact of Attractiveness on Multimodal Large Language Models

### v2

@@ Abstract
Physical attractiveness matters. It has been shown to influence human perception and decision-making, often leading to biased judgments that favor those deemed attractive in what is referred to as the “attractiveness halo effect”. While extensively studied in human judgments in a broad set of domains, including hiring, judicial sentencing or credit granting, the role that attractiveness plays in the assessments and decisions made by multimodal large language models (MLLMs) is unknown. To address this gap, we conduct an empirical study with 7 diverse open-source MLLMs evaluated on 91 socially relevant scenarios and a diverse dataset of 924 face images –corresponding to 462 individuals both with and without beauty filters applied to them. Our analysis reveals that attractiveness impacts the decisions made by MLLMs in 86.2% of the scenarios on average, demonstrating substantial bias in model behavior in what we refer to as an attractiveness bias. Similarly to humans, we find empirical evidence of the existence of the attractiveness halo effect in 94.8% of the relevant scenarios: attractive individuals are more likely to be attributed positive traits, such as intelligence or confidence, by MLLMs than unattractive individuals. Furthermore, we uncover gender, age and race biases in a significant portion of the scenarios which are also impacted by attractiveness, particularly in the case of gender, highlighting the intersectional nature of the algorithmic attractiveness bias. Our findings suggest that societal stereotypes and cultural norms intersect with perceptions of attractiveness in MLLMs in a complex manner. Our work emphasizes the need to account for intersectionality in algorithmic bias detection and mitigation efforts and underscores the challenges of addressing biases in modern MLLMs.
@@ 1 Introduction
With the advent and wide adoption of multimodal large language models MLLMs [^2] In this paper, we use the term MLLM to refer to models that integrate text and images. [Wu et al.(2023), Zhang et al.(2024b)], algorithmic biases, including the attractiveness bias and the attractiveness halo effect, have become a growing concern. Unlike large language models (LLMs) that rely solely on textual data, MLLMs are able to process both textual and visual inputs, enabling them to interpret and generate responses based on complex multimodal information. These capabilities make MLLMs highly versatile, powerful and applicable to numerous vision-and-language tasks, ranging from image captioning, visual question answering [Hu et al.(2024)] and content creation [^3] https://openai.com/index/sora-is-here/, https://www.wired.com/story/linkedin-ai-generated-influencers/?utm_source=chatgpt.com to conversational AI grounded in visual context [Pan et al.(2023), Dong et al.(2024)]. However, these models can inadvertently replicate or even magnify appearance-based biases such as the attractiveness bias. When MLLMs are trained on datasets containing images and descriptions of individuals, these models may implicitly treat attractiveness as a relevant factor, raising two key concerns: (1) MLLMs may exhibit an attractiveness bias, hence making decisions or judgments that differ based solely on an individual’s perceived attractiveness; and (2) MLLMs, like humans, may behave according to the attractiveness halo effect, thus favoring or assigning positive traits to individuals who are perceived as attractive, even when those traits are unrelated to the person’s actual abilities or character. In both cases, MLLMs could lead to a different or even preferential treatment of individuals who are considered to be more attractive by the model, potentially influencing critical decisions in areas such as hiring recommendations, educational assessments or immigration evaluations. Therefore, investigating the existence of an attractiveness bias and the presence and mechanisms of the attractiveness halo effect in MLLMs are essential to ensure fair outcomes in their deployment.
@@ 3. RQ3:
  • We design 91 scenarios and propose a methodology to examine the existence of an attractiveness bias and the attractiveness halo effect in seven distinct open-source MLLMs.
@@
The rest of the paper is organized as follows: we first provide a summary of the most relevant related research in Section 2, followed by a detailed description of the adopted methodology in Section 3. Sections 4 and 5 present our results and their discussion, respectively. Finally, our conclusions and future research directions are summarized in Section 6.
@@ 2.1 Biases in Large Language Models (LLMs)
Given the versatility of LLMs and their usefulness for a wide range of tasks, bias evaluation in LLMs has addressed a broad spectrum of scenarios and demographic groups. In fact, a variety of social biases [Yeh et al.(2023)] and biases related to reasoning and decision-making [Itzhak et al.(2024)] have been reported in LLMs, including sentiment [Huang et al.(2020)], religious [Abid et al.(2021)] and stereotype [Nadeem et al.(2021)] biases.
Gender biases in majority [Zhao et al.(2018), Vig et al.(2020), Kotek et al.(2023)] and minority [Hada et al.(2024)] languages, and disparities in the representation of various demographic groups [Narayanan Venkit et al.(2023), Lee et al.(2024), Derner et al.(2024)] have also been studied in isolation and where multiple sensitive attributes, such as gender and race, are considered simultaneously [Ma et al.(2023)]. These studies have highlighted that disparities are often more pronounced for intersectional minorities, such as Black women [Wan and Chang(2024)].
@@ 2.2 Biases in Multimodal LLMs (MLLMs)
The processing of visual information in MLLMs raises concerns about potential appearance-based biases, such as the attractiveness bias, which could influence decisions made by these models. This issue has gained increased relevance in light of emerging applications that employ MLLMs to assess attractiveness in contexts such as post-surgical evaluations [Ali and Cui(2025)], thereby amplifying the necessity of understanding how visual appearance influences model behavior.
@@ 3 Methodology
The adopted experimental methodology is summarized in Figure 1. As seen in the Figure, we formulated 91 different scenarios (described in Section 3.2) and probed the MLLMs by asking scenario-specific questions to assess the model’s attractiveness bias. In all scenarios, the input consisted of a facial image and a textual prompt containing a question about the image with two possible answers (a) or (b) that the model had to choose from.
@@ 3.1 Models
Model Model Size (# parameters)
@@ 3.2 Inputs
We leverage a dataset that was created to study the attractiveness halo effect in humans [Gulati et al.(2024)]. It consists of faces obtained from the Chicago Faces Database (CFD) [Ma et al.(2015)] and the FACES database [Ebner et al.(2010)]. The CFD provides a diverse set of face images in terms of race, with equal representation of individuals self-identifying as Asian, Black, Latino, Indian, White, or Mixed race, yet of similar ages. Conversely, the FACES database contains face images of different ages, categorized into three age groups: young, middle-aged, and old, yet with no racial diversity. As a result, the dataset comprises 462 distinct faces, with 25 images for each gender-ethnicity pair and 27 images for each gender-age group pair. All images feature individuals wearing identical clothing, displaying neutral facial expressions, and set against uniform backgrounds to minimize potential confounding factors. Furthermore, each original face image has its corresponding beautified version created using a common beautification filter available in one of the most popular selfie editing apps. Thus, for each individual in the dataset, there are the original (non-beautified) and the beautified versions, yielding a total of 924 images.
In addition to the images, the dataset contains the individual, median and mean ratings of each image provided by at least 25 human raters on a 7-point Likert scale according to the following attributes: attractiveness, intelligence, trustworthiness, sociability, happiness, femininity and unusualness. Figure 1 displays examples of the original and beautified versions of two faces from the dataset. Note that 96.1% of the faces in the dataset were rated as more attractive after applying a beauty filter and no individual was rated as less attractive after beautification. Thus, the dataset contains a diverse set of pairs of facial images where the only difference between them is attractiveness, with minimal confounds. We direct the interested reader to [Gulati et al.(2024)] for a detailed description of the dataset.
@@
The textual prompts for each of the 91 scenarios consisted of a question referring to the face image and two possible answers, indicated as (a) or (b), from which the model had to choose. Since prompt order impacts the behavior of LLMs [Lu and Yin(2021)] and MLLMs [Shi et al.(2024), Chen et al.(2024), Liu et al.(2025)], each image was evaluated under all possible orderings of the options and the average across all these orderings was used to evaluate the model, as detailed in Section 3.5. For example, a prompt could involve presenting the model with an image and asking it to determine whether the individual in the image is confident or insecure. A specific prompt for this scenario would take the form:
Make your best estimate. Is the person in the picture (a) confident or (b) insecure. Respond only with (a) or (b).
@@
Make your best estimate. Is the person in the picture (a) confident or (b) insecure. Respond only with (b) or (a).
Make your best estimate. Is the person in the picture (a) insecure or (b) confident. Respond only with (a) or (b).
Make your best estimate. Is the person in the picture (a) insecure or (b) confident. Respond only with (b) or (a).
@@ 3.3 Scenarios
Figure 2: Visual depiction of the 91 scenarios used in the experiments: 39 scenarios were about jobs, divided in 3 types, gender, race and attractiveness; 41 scenarios referred to traits; and 11 scenarios referred to conditions and were structured in 3 subgroups, geography, wealth and education/religion. The color coding reflected in the Figure is adopted in all the Figures in the remainder of the paper.
@@
Scenarios in this category involve pairs of occupations traditionally associated with specific gender (eg., “Doctor” vs “Nurse”) and racial groups (eg., “Cleaner” vs “Security guard”) or different levels of expected attractiveness (eg., “Model” vs “Makeup artist”). These scenarios aim to assess both attractiveness and societal biases rooted in stereotypes about professional roles. A total of 39 scenarios were selected in this category, informed by previous work [Cash and Kilcullen(1985), Hosoda et al.(2003)] and data from the 2022 labor force characteristics report by the U.S. Bureau of Labor Statistics (BLS) [^4] https://www.bls.gov/opub/reports/race-and-ethnicity/2022/ The full list of scenarios in this category can be found in Table 6 in Appendix C.
The scenarios in this group were divided into three types:
i) Gender-stereotyped jobs, consisting of occupations traditionally associated with one gender in Western societies, such as nursing, teaching and caregiving for women and construction, engineering and leadership roles for men. The 19 job pairs were chosen from existing datasets of stereotyped jobs [Xiao et al.(2024), Fraser and Kiritchenko(2024)], which in many cases were selected based on the latest labor force characteristics report by the U.S. Bureau of Labor Statistics [^5] https://www.bls.gov/cps/cpsaat11.htm (BLS).
@@
These scenarios feature pairs of behavioral traits, where one trait is generally perceived as positive and desirable –such as trustworthiness, confidence or friendliness– and the other as negative and undesirable –such as untrustworthiness, insecurity or hostility. The scenarios in this category were designed to evaluate the extent to which the model associates certain personal characteristics, whether favorable or unfavorable, with perceptions of attractiveness.
In total, we defined 41 scenarios that consisted of pairs of behavioral traits, selected from datasets used to analyze stereotypes in vision-language models [Hamidieh et al.(2024), Jiang et al.(2024), Zhou et al.(2022)] and are listed in Tables 7 and 8 in Appendix C. In particular, we selected a subset of behavioral traits most likely to exhibit an attractiveness bias from the 374 words proposed in the So-B-IT taxonomy [Hamidieh et al.(2024)]. To accomplish this, we created two clusters of words: one consisting of positive appearance-related terms (e.g., “attractive”, “beautiful”) and the other with negative appearance-related terms (e.g., “unattractive”, “ugly”). We then identified the top 6 words from So-B-IT that were the closest to each cluster and used GPT-4 to generate the antonyms of these words. A related methodology was employed to extract relevant word pairs from VLStereoSet [Zhou et al.(2022)]. In addition to these terms, ModSCAN [Jiang et al.(2024)] includes gender-stereotyped hobby pairs, which were also incorporated into the stereotype traits to complete the set of 41 scenarios.
@@ 3.5 Problem Formulation and Metrics
We denote with $\mathcal{X}$ the dataset of faces, where $\mathcal{X}^{o}$ is the set of original images and $\mathcal{X}^{b}$ is the set of corresponding beautified images. The dataset includes annotations for gender, age, and race for each subject, and we use a subscript to refer to specific demographic subsets. For example, the set of female faces is denoted as $\mathcal{X}_{\textit{female}}$.
@@
where $k\in\{1,\dots,K\}$ represents each of the $K$ seeds (3 in our case). For each tested scenario $s_{i}$, we collect $M\times K$ responses (4x3=12 in our case) for each image accounting for the M choice orderings and K seed combinations. To measure the number of stereotyped responses by the models, we define the Stereotype Consistency Score (SCS) as:
@@
For example, we measure the impact of attractiveness on the gender bias by evaluating the gender bias on the original (${H}_{i,\mathcal{X}^{o}}^{\textit{}gender}$) and the beautified (${H}_{i,\mathcal{X}^{b}}^{\textit{}gender}$) images for each scenario $s_{i}$ and then comparing them pairwise across all 91 scenarios using a Wilcoxon Paired Rank Test, i.e.:
@@
where WPRT denotes the Wilcoxon Paired Rank Test.
Across all tests, statistical significance and hence the existence of a bias is determined by p-values $<0.01$, and the corresponding significance levels are explicitly reported for each test. In the case of the intersectional effects, the Bonferroni correction is applied to address the multiple comparisons problem [Rupert Jr et al.(2012)]. Post the correction, the same alpha value is used as in the other tests.

### v3

@@ Abstract
Physical attractiveness matters. It has been shown to influence human perception and decision-making, often leading to biased judgments that favor those deemed attractive in what is referred to as the “attractiveness halo effect”. While extensively studied in human judgments in a broad set of domains, including hiring, judicial sentencing or credit granting, the role that attractiveness plays in the assessments and decisions made by multimodal large language models (MLLMs) is unknown. To address this gap, we conduct an empirical study with 7 diverse open-source MLLMs evaluated on 91 socially relevant scenarios and a diverse dataset of 924 face images –corresponding to 462 individuals both with and without beauty filters applied to them. Our analysis reveals that attractiveness impacts the decisions made by MLLMs in 86.2% of the scenarios on average, demonstrating substantial bias in model behavior in what we refer to as an attractiveness bias. Similarly to humans, we find empirical evidence of the existence of the attractiveness halo effect in 94.8% of the relevant scenarios: attractive individuals are more likely to be attributed positive traits, such as trustworthiness or confidence, by MLLMs than unattractive individuals. Furthermore, we uncover gender, age and race biases in a significant portion of the scenarios which are also impacted by attractiveness, particularly in the case of gender, highlighting the intersectional nature of the algorithmic attractiveness bias. Our findings suggest that societal stereotypes and cultural norms intersect with perceptions of attractiveness in MLLMs in a complex manner. Our work emphasizes the need to account for intersectionality in algorithmic bias detection and mitigation efforts and underscores the challenges of addressing biases in modern MLLMs.
@@ 1 Introduction
With the advent and wide adoption of multimodal large language models MLLMs [^2] In this paper, we use the term MLLM to refer to models that integrate text and images. [Wu et al.(2023), Zhang et al.(2024b)], algorithmic biases, including the attractiveness bias and the attractiveness halo effect, have become a growing concern. Unlike large language models (LLMs) that rely solely on textual data, MLLMs are able to process both textual and visual inputs, enabling them to interpret and generate responses based on complex multimodal information. These capabilities make MLLMs highly versatile, powerful and applicable to numerous vision-and-language tasks, ranging from image captioning, visual question answering [Hu et al.(2024)] and content creation [^3] https://openai.com/index/sora-is-here/, https://www.wired.com/story/linkedin-ai-generated-influencers/?utm_source=chatgpt.com to conversational AI grounded in visual context [Pan et al.(2023), Dong et al.(2024)]. However, these models can inadvertently replicate or even magnify appearance-based biases such as the attractiveness bias. When MLLMs are trained on datasets containing images and descriptions of individuals, these models may implicitly treat attractiveness as a relevant factor, raising two key concerns: (1) MLLMs may exhibit an attractiveness bias, hence making decisions or judgments that differ based solely on an individual’s perceived attractiveness; and (2) MLLMs, like humans, may behave according to the attractiveness halo effect, thus favoring or assigning positive traits to individuals who are perceived as attractive, even when those traits are unrelated to the person’s actual abilities or character. In both cases, MLLMs could lead to a different or even preferential treatment of individuals who are considered to be more attractive by the model, potentially influencing critical decisions in areas such as hiring recommendations or educational assessments. Therefore, investigating the existence of an attractiveness bias and the presence and mechanisms of the attractiveness halo effect in MLLMs are essential to ensure fair outcomes in their deployment.
@@ 3. RQ3:
  • We design 91 scenarios and propose a methodology to examine the existence of an attractiveness bias and the attractiveness halo effect in 7 distinct open-source MLLMs.
@@
The rest of the paper is organized as follows: we first provide a summary of the most relevant related research in Section 2, followed by a detailed description of the adopted methodology in Section 3. Sections 4 and 5 present our results and their discussion, respectively. Our conclusions and future research directions are summarized in Section 6.
@@ 2.1 Biases in Large Language Models (LLMs)
Given the versatility of LLMs and their usefulness for a wide range of tasks, bias evaluation in LLMs has addressed a broad spectrum of scenarios and demographic groups. In fact, a variety of social biases [Yeh et al.(2023)] and biases related to reasoning and decision-making [Itzhak et al.(2024)] have been reported in LLMs, including sentiment [Huang et al.(2020)], religious [Abid et al.(2021)] and stereotype [Nadeem et al.(2021)] biases. Gender biases in majority [Zhao et al.(2018), Vig et al.(2020), Kotek et al.(2023)] and minority [Hada et al.(2024)] languages, and disparities in the representation of various demographic groups [Narayanan Venkit et al.(2023), Lee et al.(2024), Derner et al.(2024)] have also been studied in isolation and where multiple sensitive attributes, such as gender and race, are considered simultaneously [Ma et al.(2023)]. These studies have highlighted that disparities are often more pronounced for intersectional minorities, such as Black women [Wan and Chang(2024)].
@@ 2.2 Biases in Multimodal LLMs (MLLMs)
The processing of visual information in MLLMs raises concerns about potential appearance-based biases, such as the attractiveness bias, which could influence decisions made by these models, both in an implicit –e.g., when evaluating candidates for positions– and explicit –e.g.. in post-surgical evaluations of plastic surgery facial procedures [Ali and Cui(2025)]– manner, thereby amplifying the necessity of understanding how visual appearance influences model behavior.
@@ 3 Methodology
The adopted experimental methodology is summarized in Figure 1. As seen in the Figure, we formulated 91 different scenarios (described in Section 3.2) and probed the MLLMs by asking scenario-specific questions to assess the model’s attractiveness bias. In all scenarios, the input consisted of a facial image and a textual prompt containing a question about the image with 2 possible answers that the model chose from.
@@ 3.1 Models
Model Name Size (# parameters)
@@ 3.2 Inputs
We leverage a dataset that was created to study the attractiveness halo effect in humans [Gulati et al.(2024)]. It consists of a curated set of faces obtained from the Chicago Faces Database (CFD) [Ma et al.(2015)] and the FACES database [Ebner et al.(2010)]. The CFD provides a diverse set of face images in terms of race, with equal representation of individuals self-identifying as Asian, Black, Latino, Indian, White, or Mixed race, yet of similar ages. Conversely, the FACES database contains face images of different ages, categorized into three age groups: young, middle-aged, and old, yet with no racial diversity. As a result, the dataset comprises 462 distinct faces, with 25 images for each gender-ethnicity pair and 27 images for each gender-age group pair. All images feature individuals wearing identical clothing, displaying neutral facial expressions, and set against uniform backgrounds to minimize potential confounds. Furthermore, for each original face image there is its corresponding beautified version created using a common beautification filter available in one of the most popular selfie editing apps. Thus, for each individual in the dataset, there are the original (non-beautified) and the beautified versions, yielding a total of 924 images.
In addition to the images, the dataset contains the self reported gender, race and age labels for the subjects in the images before the application of beauty filters. This information is considered to be the ground truth regarding the demographic information of each image. In their study with human participants, Gulati et al. [Gulati et al.(2024)] found no significant impact of the beauty filters on perceptions of gender and race by the human evaluators. The dataset also contains the individual, median and mean ratings of each image provided by at least 25 human raters on a 7-point Likert scale according to the following attributes: attractiveness, intelligence, trustworthiness, sociability, happiness, femininity and unusualness. Figure 1 displays examples of the original and beautified versions of two faces from the dataset. Note that 96.1% of the faces in the dataset were rated as more attractive after applying a beauty filter and no individual was rated as less attractive after beautification. Thus, the dataset contains a diverse set of pairs of facial images where the only difference between them is attractiveness, with minimal confounds. We direct the interested reader to [Gulati et al.(2024)] for a detailed description of the dataset.
@@
The textual prompts for each of the 91 scenarios consisted of a question referring to the face image and two possible answers, indicated as (a) or (b), from which the model had to choose. Since prompt order impacts the behavior of LLMs [Lu and Yin(2021)] and MLLMs [Shi et al.(2024), Chen et al.(2024), Liu et al.(2025)], each image was evaluated under all possible orderings of the options and the average across all these orderings was used to evaluate the model, as detailed in Section 3.5. For example, a prompt could involve presenting the model with an image and asking it to determine whether the individual in the image is confident or insecure. A specific prompt for this scenario would take the form: “Make your best estimate. Is the person in the picture (a) confident or (b) insecure. Respond only with (a) or (b)."
@@
“Make your best estimate. Is the person in the picture (a) confident or (b) insecure. Respond only with (b) or (a)."
“Make your best estimate. Is the person in the picture (a) insecure or (b) confident. Respond only with (a) or (b)."
“Make your best estimate. Is the person in the picture (a) insecure or (b) confident. Respond only with (b) or (a)."
@@ 3.3 Scenarios
Figure 2: Visual depiction of the 91 scenarios used in the experiments: 39 scenarios were about jobs, further divided in 3 types as shown in the Figure; 41 scenarios referred to traits; and 11 scenarios referred to stereotyped conditions
@@
Scenarios in this category involve pairs of occupations traditionally associated with specific gender (eg., “Doctor” vs “Nurse”) and racial groups (eg., “Cleaner” vs “Security guard”) or different levels of expected attractiveness (eg., “Model” vs “Makeup artist”). These scenarios aim to assess both attractiveness and societal biases rooted in stereotypes about professional roles. A total of 39 scenarios were selected in this category, informed by previous work [Cash and Kilcullen(1985), Hosoda et al.(2003)] and data from the 2022 labor force characteristics report by the U.S. Bureau of Labor Statistics (BLS) [^4] https://www.bls.gov/opub/reports/race-and-ethnicity/2022/. The full list of scenarios in this category can be found in Table 6 in Appendix C. The scenarios in this group were divided into three types:
i) Gender-stereotyped jobs, consisting of occupations traditionally associated with one gender in Western societies, such as nursing and caregiving for women and engineering and leadership roles for men. The 19 job pairs were chosen from existing datasets of stereotyped jobs [Xiao et al.(2024), Fraser and Kiritchenko(2024)], which in many cases were selected based on the latest labor force characteristics report by the U.S. Bureau of Labor Statistics [^5] https://www.bls.gov/cps/cpsaat11.htm (BLS).
@@
These scenarios feature pairs of behavioral traits, where one trait is generally perceived as positive and desirable –such as trustworthiness or confidence– and the other as negative and undesirable –such as insecurity or hostility. The scenarios in this category were designed to evaluate the extent to which the model associates certain personal characteristics, whether favorable or unfavorable, with perceptions of attractiveness.
In total, we defined 41 scenarios that consisted of pairs of behavioral traits, selected from datasets used to analyze stereotypes in vision-language models [Hamidieh et al.(2024), Jiang et al.(2024), Zhou et al.(2022)] and are listed in Tables 7 and 8 in Appendix C. In particular, we selected a subset of behavioral traits most likely to exhibit an attractiveness bias from the 374 words proposed in the So-B-IT taxonomy [Hamidieh et al.(2024)]. To accomplish this, we created two clusters of words: one consisting of positive appearance-related terms (e.g., “attractive”, “beautiful”) and the other with negative appearance-related terms (e.g., “unattractive”, “ugly”). We then identified the top 6 words from So-B-IT that were the closest to each cluster and used GPT-4 to generate the antonyms of these words. A related methodology was employed to extract relevant word pairs from VLStereoSet [Zhou et al.(2022)]. In addition, ModSCAN [Jiang et al.(2024)] includes gender-stereotyped hobby pairs, which were also incorporated into the stereotype traits to complete the set of 41 scenarios.
@@ 3.5 Problem Formulation and Metrics
We denote with $\mathcal{X}$ the dataset of faces, where $\mathcal{X}^{o}$ is the set of original images and $\mathcal{X}^{b}$ is the set of corresponding beautified images. Note that the dataset includes ground truth metadata corresponding to the gender, age, and race of the individual in each image.
In the following, we denote with a subscript specific demographic subsets in the dataset, e.g., the set of female faces is referred to as $\mathcal{X}_{\textit{female}}$.
@@
where $k\in\{1,\dots,K\}$ represents each of the $K$ seeds (3 in our case). For each tested scenario $s_{i}$, we collect $M\times K$ responses (4x3=12 in our case) for each image corresponding to the M choice orderings and K seed combinations. To measure the number of stereotyped responses by the models, we define the Stereotype Consistency Score (SCS) as:
@@
For example, we measure the impact of attractiveness on the gender bias by evaluating the gender bias on the original (${H}_{i,\mathcal{X}^{o}}^{\textit{}gender}$) and the beautified (${H}_{i,\mathcal{X}^{b}}^{\textit{}gender}$) images for each scenario $s_{i}$ and then comparing them pairwise across all 91 scenarios using a Wilcoxon Paired Rank Test (WPRT) as indicated below:
@@
Across all tests, statistical significance and hence the existence of a bias is determined by p-values <0.01<0.01, and the corresponding significance levels are explicitly reported for each test. In the case of the intersectional effects, the Bonferroni correction is applied to address the multiple comparisons problem [Rupert Jr et al.(2012)]. Post the correction, the same alpha value is used as in the other tests.